Thompson and Thompson is a steel bolts manufacturing company. Their current steel bolts have a mean diameter of 147 millimeters, and a variance of 36.

If a random sample of 43 steel bolts is selected, what is the probability that the sample mean would be greater than 144.9 millimeters? Round your answer to four decimal places.
Answer #1

Solution :

Given that ,

\sigma\bar x = \sigma / \sqrt n = 6 / \sqrt 43 = 0.9150

P(\bar x > 144.9) = 1 - P(\bar x < 144.9)

= 1 - P[(\bar x - \mu \bar x ) / \sigma \bar x < (144.9 - 147) / 0.9150]

= 1 - P(z < -2.30)

= 1 - 0.0107

= 0.9893

Probability = 0.9893